Two angles of polygon are 120° and 90° and the rest are each 166°. Find the number of the sides

Answer by ikleyn(52776)

.
Two angles of polygon are 120° and 90° and the rest are each 166°. Find the number of the sides
~~~~~~~~~~~~~~~~~~~~~~~~~~~~

120 + 90 + 166*(n-2) = 180*(n-2),

210 + 166n - 332 = 180n - 360,

238 = 14n,

N = 238/14 = 17 sides.

Answer. 17 sides.